OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. L(5, 0), M(0, 5)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(0-5)/(5-0)=-1

OpenStudy (alexandervonhumboldt2):

y=-x+5

OpenStudy (alexandervonhumboldt2):

do you know how to convert?

OpenStudy (anonymous):

x+y=5

OpenStudy (anonymous):

thnx